OpenStudy (anonymous):

are these the solutions of the inequality?? 2x-y<3; (0,0) (2,-2)

OpenStudy (anonymous):

Inequalities usually have plenty of solutions... use that coordinate to see if they are part of the set of solutions. for example 2(0)-0=0 which is less than 3.

OpenStudy (anonymous):

similarly 2(2)-(-2)=4+2=6 which is greater than 3 so this is not part of the set of solutions